# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ 7.2 Vision Language Model Attack
From the perspective of attack objectives, early studies pay attention to adversarial robustness tests that cause VLMs to produce incorrect descriptions (Dong et al., 2023), or evaluate the performance on out-of-distribution images — images not well represented in the training set (Tu et al., 2023). Recent work focuses more on attacks aimed at eliciting harmful content. We categorize attack approaches into three types, based on the manipulation of text, image, or cross-modal inputs.
7.2.1 Textual Prompt
@@
Small changes and design choices in the prompt can lead to significant differences in the output. The attack strategies are applicable to texts delineated in Sec. 4. For Vision-Language Models (VLMs), the methods closely mirror those used for Large Language Models (LLMs), albeit with slight modifications. For example, Maus et al. (2023) developed a black-box framework to create adversarial prompts, which can either function independently or be attached to benign prompts, to steer the generative process toward specific outcomes, such as producing images of a certain object or generating text with high perplexity.; Wu et al. (2023c) implemented four text prompt enhancement techniques: prefix injection, refusal suppression, hypothesis scenarios and appealing with emotion (belongs to psychological manipulation) to attack VLMs. Qi et al. (2023a) proposed a universal gradient-based approach that optimizes a single visual adversarial example. They initiate with a small corpus consisting of some few-shot examples of harmful content. Creation of the adversarial example is rather straightforward by maximizing the generation probability of this few-shot corpus conditioned on the adversarial example that could originate from both visual and textual input space.
Figure 18: Examples for illustrating VLMs attack methods of generating adversarial image. Example in (a) is sourced from Dong et al. (2023), (b) is from Tu et al. (2023), and (c) is from Liu et al. (2023f).
7.2.2 Adversarial Image
Vision-language tuning can weaken safety protocols planted in LLMs (Tu et al., 2023). VLMs are more vulnerable than LLMs due to the incorporation of images, thus the majority of VLMs attack methods focus on how to generate adversarial images that induce incorrect, irrelevant, or harmful text output (see example in Figure 18).
@@ 7.2.2 Adversarial Image
Dong et al. (2023) explored the adversarial robustness of Google’s Bard through the task of image captioning, where they generate adversarial image examples to elicit incorrect descriptions. Specifically, they employ white-box surrogate models to create adversarial examples via two methods: (1) an image embedding attack that causes the adversarial image’s embedding to deviate from the original image’s embedding, and (2) a text description attack aimed at maximizing the log-likelihood of a ground-truth target sentence given the adversarial image. Afterward, leveraging adversarial transferability, these examples are applied to attack commercial VLMs, showing the vulnerability of commercial visual models against adversarial attacks. [^4] Adversarial transferability assumes that adversarial examples generated for white-box models are likely to mislead black-box models.
Tu et al. (2023) generate adversarial images by maximizing the distance between the image representation and language representation of the original image description. This strategy aims to mislead a VLM into generating toxic or unrelated content.
Liu et al. (2023f) observed that when presented with a prompt-irrelevant image and a malicious question, LLaVA-1.5 demonstrates robust safety features, typically by refusing to answer or issuing warnings to users. However, when query-relevant images are used, the ASR increases significantly. They suggest that the presence of a query-relevant image triggers the model’s vision-language alignment module, which, being generally trained without safety alignment, fails to recognize harmful queries, leading to inappropriate responses. To this end, they exploit prompt-relevant images to jailbreak the open-source VLMs.
@@
(Gong et al., 2023) introduce FigStep highlighting that VLMs struggle with typographic visual prompts. They first paraphrase potentially harmful questions into step-by-step executable instructions, which are then transformed into typographic images, coupled with textual incitement to elicit dangerous outputs from models. This approach demonstrates a significantly higher ASR compared to traditional text-only attacks on open-source VLMs. However, its effectiveness diminishes in models equipped with OCR capabilities that can identify and mitigate harmful content within image prompts
@@
To mitigate the ineffectiveness of typographic attacks, Wang et al. (2023d) propose a method where the target text desired from the VLM is first converted into the target image using a T2I model. Following this, GPT4 is used to deduce a plausible instruction from the target text. A local surrogate model, which shares the visual encoder with the target model, is then employed to identify instruction-sensitive features in both the adversarial and target images. The strategy involves minimizing the difference between these feature sets to refine the adversarial image, enhancing the likelihood of achieving the intended outcome from the VLM.
7.2.3 Cross-modal Attack
He et al. (2023a) identified two factors that influence the VLM attack efficiency: inter-modality interaction and data diversity. They observe that VLMs can process image and text modalities concurrently, necessitating that attack strategies account for the interplay between these modalities. Attacks focused solely on one modality might see their effectiveness diminished by the compensatory capabilities of the other. Regarding data diversity, they note that insufficient consideration of this aspect, including text modality nuances and various image attributes like structural invariance, could impair the success of transfer attacks on VLMs. [^5] Transfer attacks refer to the situation where the adversary crafts adversarial examples on a white-box model to fool another black-box model, which is extensively applied to attack commercial models in VLM settings.
Motivated by these insights, they proposed a dual-pronged approach to attacks, incorporating both text and image dimensions. The text component is to replace vulnerable words, while the image aspect seeks adversarial images that diverge maximally from the associated benign text in a high-dimensional manifold.
Shayegani et al. (2023a) delve into the vulnerabilities related to cross-modal alignment with their pioneering study on compositional adversarial attacks within the aligned embedding space. They craft four distinct scenarios, each pairing a harmless text instruction with a malevolent image, to dissect and understand the nuances of harmful prompt decomposition.
@@ 7.3 Benchmarks
There are a few benchmarks focusing on vision-language model safety evaluation.
Safebench (Gong et al., 2023) is constructed by first collecting 10 topics prohibited by OpenAI and Meta policies, and then prompting GPT4 to generate 50 questions under each topic, followed by human review, resulting in 500 questions.
Query-relevant Text-image Pair: benchmark, devised by (Liu et al., 2023f), involves a four-step creation process. This process includes generating malicious questions across 13 harmful categories using GPT4, extracting unsafe phrases, converting these queries into images using a stable diffusion model and typography tools, and finally, rephrasing the questions. This results in a comprehensive dataset of 5040 samples.
RTVLM is the first red teaming dataset to benchmark current VLMs in terms of four primary aspects: faithfulness, privacy, safety, and fairness, with a total of 5,200 (image, question, refusal flag, reference answer) pairs (Li et al., 2024c). It encompasses 10 subtasks such as text/image misleading, multimodal attacking, and face fairness. Unlike other VLM benchmarks with only (image and text) pairs, each RTVLM data instance is composed of two additional parts: (1) a label indicating whether the question is safe to answer, and (2) reference answers generated by humans or GPT4. These two features serve as the “golden standard” in the evaluation process.
@@ 7.4 Safeguards
Unlike text which consists of discrete words, multimodal information usually includes images, videos, audio, etc., which have two dimensional structure, and unlike text, they are usually continuous in the input space. In contrast to text-based generative models, multi-modal large models are susceptible to malicious attacks on visual inputs, posing significant challenges to alignment mechanisms due to the continuity of image signals.
Pi et al. (2024) proposed MLLM-Protector, comprising a harm detector and a lightweight classifier that evaluates the harmfulness of responses generated by the VLM. If the output is deemed potentially harmful, a response detoxifier is activated to adjust the output, ensuring compliance with safety standards. This approach effectively reduces the risk of harmful outputs without compromising the overall performance of VLM.
@@ 8 LLM-based Application Red Teaming
As large language models are widely used for various applications and scenarios, the problem of how to perform red teaming for LLM-based applications has attracted the attention of the research community. Some works focus on red teaming LLM agent systems. In this survey, we use application to refer to the whole system and LLM to refer to the large language model they rely on. We do not differentiate applications and LLM agent systems.
@@
Program Program Development Terminal, Code Edit, Github, Code Security
@@
8.1 Application Scenarios and Risks
With the powerful instruction-following capability, LLMs are endowed with diverse tools and interfaces to perform real-life tasks, including general chat, browsing web pages, making orders, etc. This enables them to be used in a variety of scenarios. Yuan et al. (2024) categorized the application scenarios into 7 categories, as listed in Table 3.
@@ 8.1 Application Scenarios and Risks
For different scenarios, applications may adopt different system designs. This includes the system prompts for LLMs, the callable tools, and how the information flows among components in applications. Many studies have shown the vulnerability of LLMs, which may lead to different types of risks. However, new risks are raised because of the security flaws in these scenarios. For example, when LLMs have access to the bank account or can even operate it, there is the risk of financial loss, which is scarce for raw LLMs. Therefore, risks for LLMs and applications are different.
@@
Current works focus mainly on the following types of risks for LLM-based applications: privacy leakage (breach), financial loss, inaccurate (inefficient) execution, safety hazards, physical harm, reputation damage, computer security, illegal activities, data loss, property damage, ethic, and moral hazards, bias and offensiveness, and other miscellaneous risks (Ruan et al., 2023; Yuan et al., 2024). There are also some works focus on the application in a specific domain, e.g.: scientific domains (Tang et al., 2024).
8.2 Attack Methods
@@
LLM-based applications inherit the vulnerability of their foundation model. Therefore, methods for attacking LLMs may also be applied to attack these applications. Besides, endowing LLMs with different capabilities also introduces new vulnerabilities and new attacking methods aim to exploit them. In this section, we first categorize the application-specific attack methods into LLM-based and application-based methods, respectively. Then we explain the vulnerabilities of applications.
LLM-based Attacking
@@
This type of work tries to attack the LLMs that the application is based on and hence is discussed in a dedicated section Sec. 4. Current applications usually assign LLMs with different roles by setting system prompts according to their use cases. Therefore, some works explore re-configuring system prompts to make the applications generate malicious content (Tian et al., 2023b; Zhang et al., 2024e; b). A typical way is to inject negative personalities, which largely alter LLMs’ behaviors and make them generate harmful content. Some works deliberately add backdoors (Yang et al., 2024) in system prompts, which can be triggered by user queries during later conversations. As LLMs in applications are usually customized to communicate with users, some works attack them directly with malicious user queries (Ye et al., 2024; Zhang et al., 2024e) by requesting LLM-based applications to generate harmful content using jailbreak techniques.
Application-based Attacking
@@
These works attack applications based on the nature that LLMs need to interact with environments to obtain results or observations after they have executed some actions. One line of work poisons the observations returned by environments or tools. Deng et al. (2024) integrated malicious content into documents that might be retrieved by LLMs. Yang et al. (2024) injected backdoor triggers in observations. These documents and observations are input into LLMs, causing them to generate harmful content. Some other works attack applications by attacking their integrated tools or interfaces. Because tools and interfaces are often application-specific, these methods are often application-specific. Pedro et al. (2023) made LLMs generate unsafe SQL queries and execute them, causing unexpected results such as data loss and unauthorized data access.
@@
Many works have shown that LLM-based application is more vulnerable than bare LLMs (Tian et al., 2023b; Yu et al., 2023b). One of the reasons is the misalignment between LLM’s concepts of safety and the constraints in application scenarios. For example, an LLM may consider generating an uninspected deletion SQL query as a safe behavior while this may cause catastrophic data loss if it is executed by applications. Also, (Antebi et al., 2024) show that OpenAI GPTs can be attacked with various methods. Multi-agent systems are shown to be more vulnerable than single-agent systems. Jailbreak for one LLM can cause the whole system to be affected (Gu et al., 2024). Also, for a hierarchical agent system, when high-level LLMs (e.g. planners) are jailbroken, they tend to induce low-level LLMs (e.g. action executers) to generate harmful content (Tian et al., 2023b).
8.3 Defense
Since many attacking methods aim to attack the LLMs of applications, defending strategies can also be divided into LLM-based and application-based.
LLM-based Defense
This line of work focuses on preventing LLMs from generating malicious content. Techniques that are used to defend LLMs can also be applied (refer to Sec. 6). However, Zhang et al. (2024e) argue that simply conducting defense at the input stage is insufficient. They further propose psychological assessment defense before tool executions and role defense by designing a police role. Additionally, since LLMs need to interact with environments or execute tools, a major difference between defending raw LLMs and applications is using risk awareness. Risk awareness indicates that LLMs know possible risk types and risk descriptions in current scenarios before action. Several works demonstrated that risk awareness is helpful for LLMs to behave safely (Ruan et al., 2023; Yuan et al., 2024). However, these possible risks may not often be accessible. Instead, (Hua et al., 2024) retrieve several safety regulations and take them as part of input.
Application-based Defense
Although LLMs may be safer with the above strategies, attacking can still be executed on integrated components in applications. For methods designed to attack tools, the defense can be implemented by checking the query used to call tools or customizing tool permissions (Pedro et al., 2023).
8.4 Evaluation

### v2

@@ 7.2 Vision Language Model Attacks
From the perspective of attack objectives, early work paid attention to adversarial robustness tests that cause VLMs to produce incorrect descriptions Dong et al. (2023), or evaluate the performance on out-of-distribution images, i.e., images not well represented in the training set Tu et al. (2023). Recent work has focused more on attacks aimed at eliciting harmful content. We categorize attack approaches into three types, based on whether they manipulate the text, image, or cross-modal inputs.
7.2.1 Textual Prompt Attacks
@@
Small changes and design choices in the prompt can lead to significant differences in the output. The attack strategies targeted at text, as described in Section 4, are largely applicable to Vision-Language Models (VLMs), with minor modifications. For example, Maus et al. (2023) developed a black-box framework to create adversarial prompts, which can either function independently or be attached to benign prompts, to steer the generative process toward specific outcomes, such as producing images of a certain object or generating text with high perplexity. Wu et al. (2023d) implemented four text prompt enhancement techniques: prefix injection, refusal suppression, hypothesis scenarios, and appealing with emotion as forms of psychological manipulation to attack VLMs. Qi et al. (2024) proposed a universal gradient-based approach that optimizes a single visual adversarial example. They initiate with a small corpus consisting of few-shot examples of harmful content. Adversarial examples are generated simply by maximizing the generation probability of this few-shot corpus conditioned on the adversarial example that could originate from both visual and textual input space.
7.2.2 Adversarial Image Attacks
Vision-language tuning can weaken safety protocols embedded in LLMs Tu et al. (2023). VLMs are more vulnerable to attacks than LLMs due to the more unstructured nature of images compared to text, opening up possibilities both in adding adversarial noise into images to generate inappropriate text, as well as generating inappropriate textual content within images. See Figure 28 for examples.
Figure 28: Examples of VLM attack methods for generating adversarial images: (a) The model refuses to describe the original image but provides a detailed description of the perturbed image. (b) A benign instruction paired with an unrelated adversarial image prompts the model to produce harmful speech. Example in (a) is sourced from Dong et al. (2023), (b) is from Tu et al. (2023).
@@ 7.2.2 Adversarial Image Attacks
Dong et al. (2023) explored the adversarial robustness of Google’s Bard through the task of image captioning, where they generate adversarial image examples to elicit incorrect descriptions. Specifically, they employ white-box surrogate models to create adversarial examples via two methods: (1) an image embedding attack that causes the adversarial image’s embedding to deviate from the original image’s embedding, and (2) a text description attack aimed at maximizing the log-likelihood of a ground-truth target sentence given the adversarial image. Next, leveraging adversarial transferability, these examples are applied to attack commercial VLMs, showing the vulnerability of commercial visual models against adversarial attacks. [^5] Adversarial transferability assumes that adversarial examples generated for white-box models are likely to mislead black-box models.
Tu et al. (2023) generate adversarial images by maximizing the distance between the image representation and language representation of the original image description. This strategy aims to mislead a VLMs into generating toxic or unrelated content.
Liu et al. (2023f) observed that when presented with a prompt-irrelevant image and a malicious question, LLaVA-1.5 demonstrates robust safety features, typically refusing to answer or issuing warnings to the user. However, when query-relevant images are used, the ASR increases significantly. They suggest that the presence of a query-relevant image triggers the model’s vision-language alignment module, which, being trained without safety alignment, fails to recognize harmful queries, leading to inappropriate responses. To this end, they exploit prompt-relevant images to jailbreak the open-source VLMs.
@@
Gong et al. (2023) introduced FigStep, highlighting that VLMs struggle with typographic visual prompts. They first paraphrase potentially harmful questions into step-by-step executable instructions, which are then transformed into typographic images, coupled with textual incitement to elicit dangerous outputs from models. This approach demonstrates a significantly higher ASR compared to traditional text-only attacks on open-source VLMs. However, its effectiveness is diminished in models equipped with OCR capabilities that can identify and mitigate harmful content within image prompts.
@@
To mitigate the ineffectiveness of typographic attacks, Wang et al. (2023d) propose a method where the target text desired from the VLMs is first converted into the target image using a T2I model. Following this, GPT4 is used to deduce a plausible instruction from the target text. A local surrogate model, which shares the visual encoder with the target model, is then employed to identify instruction-sensitive features in both the adversarial and target images. The strategy involves minimizing the difference between these feature sets to refine the adversarial image, enhancing the likelihood of achieving the intended outcome from the VLMs.
7.2.3 Cross-modal Attacks
He et al. (2023a) identified two factors that influence VLM attack efficiency: inter-modality interaction and data diversity. They observe that VLMs can process image and text modalities concurrently, necessitating that attack strategies account for the interplay between these modalities. Attacks focused solely on one modality might see their effectiveness diminished by the compensatory capabilities of the other. Regarding data diversity, they note that insufficient consideration of this aspect, including text modality nuances and various image attributes like structural invariance, could impair the success of transfer attacks on VLMs. [^6] Transfer attacks refer to the situation where the adversary crafts adversarial examples on a white-box model to fool another black-box model, which is extensively applied to attack commercial models in VLM settings.
Motivated by these insights, they proposed a dual-pronged approach to attacks, incorporating both text and image dimensions. The algorithm iteratively optimizes a set of modified text and images to produce outputs contradictory to the original labels, starting from benign inputs. It first generates an intermediate adversarial text representation using the benign text and images, then uses this text to craft adversarial image representations, and finally refines both text and image adversarial representations in tandem. This iterative process establishes a strong connection between adversarial representations across the two modalities.
Shayegani et al. (2023a) delve into the vulnerabilities related to cross-modal alignment with their pioneering study on compositional adversarial attacks within the aligned embedding space. They craft four distinct scenarios, each pairing a harmless text instruction with a malicious image, to dissect and understand the nuances of harmful prompt decomposition.
Their findings indicate that attack success rates increase when benign textual instructions align at the embedding level with malicious triggers embedded in the visual modality. This underscores significant cross-alignment vulnerabilities in multi-modal models.
@@ 7.3 Benchmarks
There are a few benchmarks focusing specifically on vision-language model safety evaluation.
Safebench Gong et al. (2023) was constructed by first collecting 10 topics prohibited by OpenAI and Meta policies, and then prompting GPT4 to generate 50 questions under each topic, followed by human review, resulting in 500 questions.
Query-relevant Text-image Pair benchmark, devised by Liu et al. (2023f), involves a four-step creation process. This process includes generating malicious questions across 13 harmful categories using GPT4, extracting unsafe phrases, converting these queries into images using a stable diffusion model and typography tools, and finally, rephrasing the questions. This results in a comprehensive dataset of 5040 samples.
RTVLM is the first red teaming dataset to benchmark current VLMs in terms of four primary aspects: faithfulness, privacy, safety, and fairness, with a total of 5,200 (image, question, refusal flag, reference answer) pairs Li et al. (2024d). It encompasses 10 subtasks such as image misleading that investigate the models’ ability to generate accurate outputs despite given misleading inputs, multimodal attacks, and face fairness. Unlike other VLM benchmarks with only image–text pairs, each RTVLM data instance is composed of two additional parts: (1) a label indicating whether the question is safe to answer, and (2) reference answers generated by humans or GPT4.
@@ 7.4 Safeguards
In contrast to text-based generative models, multimodal large models are susceptible to malicious attacks on visual inputs, posing significant challenges to alignment mechanisms due to the continuity of image signals. Pi et al. (2024) proposed MLLM-Protector, comprising a harm detector and a lightweight classifier that evaluates the harmfulness of responses generated by VLMs. If the output is deemed potentially harmful, a response detoxifier is activated to adjust the output, ensuring compliance with safety standards. This approach effectively reduces the risk of harmful outputs without compromising the overall performance of VLMs.
@@ 8 LLM-based Application Red Teaming
Due to their impressive capabilities and versatility, LLMs are increasingly being used as agents in various real-world applications to interact with external environments and systems, in what are called LLM-Integrated Applications Liu et al. (2023i). The safety of such applications is much more complex and impactful than traditional chat-oriented safeguards, as they can result in direct and lasting impacts on individuals and the world (such as financial transfers, legally-binding transactions, or direct manipulation of the physical world). LLMs being manipulated to buy expensive plane tickets, send inappropriate emails, or bias the evaluation of individual resumes or project reports have real-world consequences.
In this section, we focus on the risks, attack methods, and defense strategies arising from integrating application-oriented techniques, such as tool calling Ye et al. (2024); Wang et al. (2024c), ReACT Yao et al. (2023c), reflection mechanisms Shinn et al. (2023), and multi-LLM communication Du et al. (2023a).
8.1 Application Scenarios and Risks
Due to their powerful instruction-following capability, LLMs can be used to interact with diverse tools and interfaces to perform real-life tasks, including writing and sending an email, placing orders, or performing a financial transaction. This enables them to be used in a variety of scenarios, which Yuan et al. (2024) categorized into 7 categories, as listed in Table 5.
@@ 8.1 Application Scenarios and Risks
Program Program Development Terminal, Code Editing, GitHub, Code Security
@@
For different scenarios, applications may adopt different system designs. This includes the style of system prompt, the callable tools, and how the information flows among components in applications. In addition to the vulnerabilities documented above for LLMs, new risks arise because of security flaws in these scenarios. For example, when LLMs have access to a bank account, there is the risk of financial loss, which is not the case for “pure” LLMs.
@@
Current work focuses mainly on the following types of risks for LLM-based applications: privacy leakage, financial loss, inaccurate (or inefficient) execution, safety hazards, physical harm, reputation damage, computer security, illegal activities, data loss, property damage, ethic and moral hazards, bias and offensiveness, and other miscellaneous risks Ruan et al. (2023); Yuan et al. (2024). There is also much work that focuses on risks in specific domains, e.g., scientific domains Tang et al. (2024).
8.2 Attacking LLM Applications
@@
LLM-based applications inherit the vulnerability of their underlying foundation model. Therefore, methods for attacking LLMs may also be applied to attack these applications (see Section 4).
@@
A notable body of work on attacking LLM applications is based on assigning the LLM with different roles by setting system prompts according to their use cases. Some work explores re-configuring system prompts to make the applications generate malicious content Tian et al. (2023b); Zhang et al. (2024e; b). A typical way is to inject negative personalities, to alter LLM behavior and make them respond negatively. Some work has deliberately added backdoors Yang et al. (2024b) in system prompts, which can be triggered by user queries during later interactions.
@@
Other work has exploited new vulnerabilities to attack the applications (rather than the LLMs), based on the fact that LLMs need to interact with an environment to obtain results or observations after they have executed some actions. One line of work poisons the observations returned by environments or tools. Deng et al. (2024) integrated malicious content into documents that might be retrieved by LLMs. Yang et al. (2024b) injected backdoor triggers into observations. These documents and observations are fed into LLMs, causing them to return malicious results. Some other work has attacked applications by attacking their integrated tools or interfaces. Because tools and interfaces are often application-specific, these methods are often application-specific. Pedro et al. (2023) made LLMs generate unsafe SQL queries to be executed, causing unexpected results such as data loss or unauthorized data access.
@@
Much work has shown that LLM-based applications are more vulnerable than pure LLMs Tian et al. (2023b); Yu et al. (2023b). One reason is the misalignment between the LLM’s concepts of safety and the constraints in application scenarios. For example, an LLM may consider generating an uninspected deletion SQL query as a safe behavior while this may cause catastrophic data loss if it is executed by an application. Also, Antebi et al. (2024) show that OpenAI GPTs can be attacked with various methods. Multi-agent systems have been shown to be more vulnerable than single-agent systems. Akin to a ripple effect, jailbreaking one LLM can cause the whole system to be affected Gu et al. (2024). Also, for a hierarchical agent system, when high-level LLMs (e.g., planners) are jailbroken, they tend to induce low-level LLMs (e.g., action executers) to act in an unsafe way Tian et al. (2023b).
8.3 Defense of LLM Applications
Many attack methods target the LLMs within applications, so standard LLM defense techniques (see Section 5) can often be applied. Building upon these, Zhang et al. (2024e) propose additional measures such as psychological assessment defense before tool execution and role-based defense by assigning a “police” role. Psychological assessment defense involves evaluating the mental state of the LLM to detect any signs of deviation from its intended behavior before it performs any actions. This can be particularly effective as it targets the internal state of the model, which may be influenced by dark personality traits injected by attackers.
Role-based defense, another strategy, is especially effective in multi-agent systems. It involves designating specific agents to act as overseers or “policemen”. These agents are tasked with monitoring the interactions and outputs of other agents within the system. They intervene when they detect behavior that deviates from acceptable norms, challenging and correcting actions that could lead to harmful outcomes. This approach is beneficial as it harnesses the collective intelligence of the multi-agent system to enforce safety standards and maintain the integrity of the system’s operations.
A key difference in defending raw LLMs versus applications is the need for risk awareness. Risk awareness allows the LLM to understand potential risks in a scenario before taking action, which has been shown to enhance safety Ruan et al. (2023); Yuan et al. (2024). Since these risks are not always readily identifiable, Hua et al. (2024) suggest retrieving safety regulations and incorporating them into the input.
While these strategies can improve safety, there is still the risk of attacks on integrated components in applications. For tool-related attacks, defense mechanisms include validating tool query inputs and customizing tool permissions Pedro et al. (2023).
8.4 LLM Applications Safety Evaluation